# Repository-level Code Translation Benchmark Targeting Rust

### v3

@@ 4.2.2. Results
The results reveal that RustRepoTrans effectively challenges LLMs in complex code translation, as even the best model, Claude-3.5, achieves a Pass@1 of only 45.3%. After one round of debugging, its DSR@1 increases to 56.5%, showing limited improvement.
@@ 4.3. RQ2: RustRepoTrans Effectiveness
To investigate the effectiveness of our newly constructed benchmark, we conducted a comparative analysis based on prior literature.
@@
To provide a comprehensive comparison, we specifically analyzed the model’s performance on both RustRepoTrans and the previously established benchmarks. This analysis aims to demonstrate how our dataset serves as a more robust test of model capabilities in complex, real-world code translation scenarios.
Benchmark Selection. Our analysis is grounded in the work of Pan et al. (Pan et al., 2024), who conducted a thorough assessment of existing code translation benchmarks. They evaluated the performance of several prominent datasets, enabling us to contextualize our findings within the broader landscape of code translation research. Consequently, we compare our benchmark with datasets such as CodeNet (Puri et al., 2021), Avatar (Ahmad et al., 2021), EvalPlus (Liu et al., 2024), as shown in Table 3.
Model Selection. For our analysis, we selected GPT-4 due to its consistent ranking as the top-performing model across various datasets in prior studies (Pan et al., 2024). While Claude-3.5 has emerged as a leading model in RustRepoTrans, it was excluded from our comparison because it is a recent development and was not included in the earlier assessments.
@@ 4.3.2. Results
The results presented in Table 3 reveal significant differences in the Pass@1 scores of GPT-4 across various datasets. Notably, while GPT-4 achieves high accuracy rates—ranging from 52.2% to 85.5%—on established benchmarks such as CodeNet, Avatar, and EvalPlus, its performance drastically drops when evaluated on RustRepoTrans. The translation results of GPT-4 on datasets other than RustRepoTrans are directly taken from the findings reported in (Pan et al., 2024).
Table 3. Pass@1 of GPT-4 in translating code from different studied datasets
dataset
Source
Language
Target Language Pass@1 (GPT-4)
CodeNet(Puri et al., 2021) C C++, Go, Java, Python 83.0%
C++ C, Go, Java, Python 80.0%
Go C, C++, Java, Python 85.5%
Java C, C++, Go, Python 81.3%
Python C, C++, Go, Java 79.9%
Avatar(Ahmad et al., 2021) Java C, C++, Go, Python 70.8%
Python C, C++, Go, Java 52.2%
EvalPlus(Liu et al., 2024) Python Java 79.3%
RustRepoTrans C Rust 29.5%
Java 31.5%
Python 27.6%
Specifically, the Pass@1 scores for RustRepoTrans are significantly lower, with rates of 29.5% for C to Rust, 31.5% for Java to Rust, and 27.6% for Python to Rust translations, averaging 29.3%. This indicates a decline of approximately 41.5% to 56.2% compared to the higher scores observed on other benchmarks. This notable decrease highlights the increased complexity and unique challenges posed by RustRepoTrans, reinforcing the need for models to effectively manage dependencies and navigate intricate code structures in real-world. While GPT-4 is proficient in translating code across more straightforward datasets, it struggles significantly when faced with the complexities inherent in translating to Rust, further validating the robustness of our benchmark in assessing model capabilities.
Additionally, Rust, as a programming language, poses additional challenges due to its stricter syntax requirements and less abundant resources during LLM training compared to more popular languages like Python and Java. This results in a higher level of difficulty when translating to Rust, further validating the robustness of our benchmark in assessing model capabilities in tackling these complex coding tasks.
@@ 4.3.3. Summary
Our comparison reveals that our benchmark significantly increases complexity and dependency management, resulting in a notable decrease in Pass@1 scores for GPT-4, from 85.5% to 29.3%. These findings indicate that our benchmark provides a more effective measure of model performance in translating intricate code structures and highlights the necessity for models that can better manage real-world project dependencies.
@@ 4.4.2. Results
Figure 8. Distribution of Successful and Unsuccessful Tasks with Claude Based on Dependencies and Source/Target Code Lines
@@ 4.5.2. Results
Table 4. The comparison of the proportions of error types between RustRepoTrans and (Pan et al., 2024).
Source Language Target Language Compilation Errors Runtime Errors Functional Errors Non-terminating Execution
@@
RustRepoTrans C++, Java, Python Rust 94.8% 0% 5.2% 0%
@@
E0599 4,601 This error occurs when a method is used on a type which doesn’t implement it.
@@
E0277 2,540 You tried to use a type which doesn’t implement some trait in a place which expected that trait.
@@
E0061 222 An invalid number of arguments was passed when calling a function.
@@
E0282 79 The compiler could not infer a type and asked for a type annotation.
@@
The details of the 10 error causes from the types are introduced as following.
@@
Comparison of Error Causes Distribution across Different LLMs. Fig. 11 shows the distribution of error causes across various LLMs during code translation on RustRepoTrans. The Failing to Understand Target Language Features error causes are relatively evenly distributed among LLMs due to their similar training data, primarily derived from publicly available code. This results in comparable understanding of the target language across models. In contrast, Failing to Understand Differences Between Languages and Other error causes are more prevalent in LLMs with weaker translation capabilities. The former assesses the model’s grasp of programming language differences, while the latter reflects its ability to follow instructions, both of which are positively correlated with translation performance.
@@
Figure 11. Comparative Distribution of Error Causes in LLMs
@@ 4.6. RQ5: Key Capabilities
To understand LLMs’ code translation Capabilities beyond overall success rates (Pass@1 in RQ1), we analyze three aspects: the Capability of noise robustness, syntactical difference identification, and code simplicity.
@@ 4.6.1. Design
The key abilities for these three aspects were tested by evaluating each model on a dataset extracted or constructed from RustRepoTrans, tailored to the specific capability being assessed. The same translation prompts used in 4.2 were applied. The experimental design for each aspect is described as follows.
@@
Figure 12. Example of checks related to language syntactic features to evaluate LLM’s capability to identify syntactical differences
@@ 4.6.2. Results
In terms of CC Rate, Claude-3.5 again excels with a CC Rate of 103.7%, while Llama-3.1-8B is notably lower at 53.6%. This difference arises because Claude-3.5 often uses concise built-in functions in place of loops. For instance, as illustrated in Fig. 15, when tasked with finding the first element in an array that meets a condition, Claude-3.5 employs the target language’s any method from the iterator library, whereas other LLMs rely on traditional loops. Furthermore, Claude-3.5’s CC Rate exceeding 1 indicates that, in terms of cyclomatic complexity, its generated code is simpler than the reference code. Fig. 15 provides an example where C serves as the source programming language. This simplification may be attributed to the Rust ground truth developers using the C2Rust migration tool (C2Rust, 2024), which translates code via one-to-one pattern matching.
@@ 6. Conclusion
7. Data Availability
All data/code used in this study is provided in the replication package (rep, 2024).

### v4

@@ 4.2.2. Results
The results reveal that RustRepoTrans effectively challenges LLMs in complex code translation, as even the best model, Claude-3.5, only achieves a Pass@1 of 45.3%, and DSR@1 of 56.5%, highlighting the limitations of LLMs.
@@ 4.3. RQ2: RustRepoTrans Effectiveness
To assess our benchmark’s effectiveness, a comparative analysis was conducted based on prior literature.
@@
We evaluated performance on RustRepoTrans and a prior Rust translation benchmark without repository-level context, demonstrating RustRepoTrans’s effectiveness in assessing models for complex, real-world translation scenarios. Our analysis is based on CodeTransOcean (Yan et al., 2023), which includes translation tasks across 45 programming languages, including Rust, without repository-level dependencies. To ensure a fair and efficient comparison, we constructed a subset of CodeTransOcean by selecting translation pairs where Rust is the target language and C, Java, or Python is the source language, removing those with non-compilable Rust code. From the remaining pairs, we randomly sampled 300 (100 per language) to match our dataset’s scale for comparative experiments.
Claude-3.5 was selected for its top performance in RustRepoTrans. It generated one translation per task, evaluated using corresponding tests. We computed the average Pass@1 across source languages and compared it with RQ1. Prompts followed those in Fig. 3 but excluded dependency-related instructions. Hyperparameters were the same as in RQ1, with the temperature set to zero.
@@ 4.3.2. Results
The results presented in Table 3 reveal significant differences in the Pass@1 scores of Claude-3.5 across various datasets. Notably, while Claude-3.5 achieves high accuracy rates—ranging from 67.0% to 80.0%—on established benchmarks, its performance drastically drops when evaluated on RustRepoTrans.
Specifically, the Pass@1 scores for RustRepoTrans are significantly lower, with rates of 48.3% for C to Rust, 39.3% for Java to Rust, and 41.7% for Python to Rust translations, averaging 42.5%. This indicates a decline of approximately 25.3% to 36.7% compared to the higher scores observed on CodeTransOcean. This notable decrease highlights the increased complexity and unique challenges posed by RustRepoTrans, reinforcing the need for models to effectively manage repository-level dependencies and navigate intricate code structures in real-world.
Table 3. Pass@1 Comparison of Claude-3.5 in C, Java, and Python to Rust on CodeTransOcean and RustRepoTrans
Source Language Dataset Pass@1(Claude-3.5)
C CodeTransOcean 80.0%
RustRepoTrans 48.3%(↓31.7%)
Java CodeTransOcean 76.0%
RustRepoTrans 39.3%(↓36.7%)
Python CodeTransOcean 67.0%
RustRepoTrans 41.7%(↓25.3%)
@@ 4.3.3. Summary
Our benchmark’s repository-level context poses greater challenges for LLMs, reducing Claude-3.5’s Pass@1 from 74.3% to 43.5% (a 30.8% drop) in average. This underscores the benchmark’s effectiveness in evaluating real-world code translation and the need for models to better handle project dependencies.
@@ 4.4.2. Results
Figure 5. Success and Failure Distribution of Claude by Dependencies and Source/Target Code Lines
@@ 4.5.2. Results
Table 4. The comparison of the proportions of error types between RustRepoTrans and (Pan et al., 2024).CE: Compilation Errors, RE: Runtime Errors, FE: Functional Errors, NTE: Non-terminating Execution
Source Target CE RE FE NTE
@@
RustRepoTrans C, Java, Python Rust 94.8% 0% 5.2% 0%
@@
E0599 4,601
This error occurs when a method is used on a type
which doesn’t implement it.
@@
E0277 2,540
You tried to use a type which doesn’t implement
some trait in a place which expected that trait.
@@
E0061 222
An invalid number of arguments was passed
when calling a function.
@@
E0282 79
The compiler could not infer a type and asked for
a type annotation.
@@
The detailed 10 error causes from the types are as following.
@@
Figure 8. Comparative Distribution of Error Causes in LLMs
Comparison of Error Causes Distribution across Different LLMs. Fig. 8 shows the distribution of error causes across various LLMs during code translation on RustRepoTrans. The Failing to Understand Target Language Features error causes are relatively evenly distributed among LLMs due to their similar training data, primarily derived from publicly available code. This results in comparable understanding of the target language across models. In contrast, Failing to Understand Differences Between Languages and Other errors are more common in LLMs with weaker translation capabilities. The former evaluates the model’s understanding of language differences, while the latter reflects its ability to follow instructions, both positively correlating with translation performance.
@@ 4.6. RQ5: Key Capabilities
Beyond Pass@1 in RQ1, we analyze LLMs’ translation capabilities in noise robustness, syntactical difference identification, and code simplicity.
@@ 4.6.1. Design
The key abilities for these three aspects were tested by evaluating each model on a dataset extracted or constructed from RustRepoTrans, tailored to the specific capability being assessed. The same translation prompts used in Section 4.2 were applied. The experimental design for each aspect is described as follows.
@@
Figure 9. Examples of checks from other languages that are not needed in Rust
@@ 4.6.2. Results
In terms of CC Rate, Claude-3.5 again excels with a CC Rate of 103.7%, while Llama-3.1-8B is notably lower at 53.6%. This difference arises because Claude-3.5 often uses concise built-in functions in place of loops. For instance, as illustrated in Fig. 12, when tasked with finding the first element in an array that meets a condition, Claude-3.5 employs the target language’s any method from the iterator library, whereas other LLMs rely on traditional loops. Furthermore, Claude-3.5’s CC Rate exceeding 1 indicates that, in terms of cyclomatic complexity, its generated code is simpler than the reference code. Fig. 12 shows an example with C as the source language. This simplification likely results from Rust ground truth developers using the C2Rust migration tool (C2Rust, 2024), which translates code through one-to-one pattern matching.